OpenStudy (anonymous):

3sqrtx^3 / 5sqrtx^2 I am really confussed.

OpenStudy (anonymous):

\[\frac{ 3\sqrt{x^{3}} }{ 5\sqrt{x^{2}} }\] is this what the question looks like ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

and you want to simplify it ?

OpenStudy (anonymous):

yes that is correct.

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 } \sqrt{\frac{ x^{3} }{ x^{2} }} = \frac{ 3 }{ 5 } \sqrt{x^{3}x^{-2}}\] I hope you can continue from here.

OpenStudy (anonymous):

yes I think I can. Thank you very much. I was not sure were to start.

OpenStudy (anonymous):

glad to help :)

OpenStudy (anonymous):

3/5 sqrt x3-2 = 3/5 sqrt x????

OpenStudy (schrodinger):

It should. As long as you have sqrt(x) in the numerator, like \[\frac{ 3\sqrt{x} }{ 5 }\]I think you should be fine. Just as a cautionary note, i'm prone to make stupid mistakes ("It's just simple maths, right?") so I might be wrong somehow, although everything I know points otherwise.